averyyy:

Zoe keeps track of the miles per gallon her car gets per week. She has accumulated the following data: (1, 24), (2, 24.38), (3, 24.76), (4, 25.14) What is the common difference or ratio? (1 point) The common difference is 0.38. The common difference is 0.76. The common ratio is 0.99. The common ratio is 1.05.

generation2006:

In order to find the common difference of the data in the problems you have to follow those steps. We know that the arithmetic sequence the difference between one term and the next is a constant in called common difference. Let 1x=1 y1=24 2x=2y 2=24.38 3x=3y 3=24.76 4x=4y 4=25.15 so y2-y1=24.38-24=0.38 y3-y2=24.76-24=0.38 y4-y3=25.15-24=0.38 the difference of the is the result of let would be the correct answer so in a simpler way to explain it subtract 24 from 24.38 and that would be the answer